OpenStudy (ttop0816):

please help! (: Solve sin2x + 2sinx = 0 for 0 ≤ x < 2 π .

OpenStudy (ttop0816):

{0} {0, π/2} {0, π} <<<< {π/2,3π/2 }

OpenStudy (ttop0816):

@kc_kennylau

OpenStudy (anonymous):

Simple

OpenStudy (anonymous):

Now here we will have 2 equation cosx +1 =0 cosx=-1 x=pi sinx=0 x=pi or 0

OpenStudy (anonymous):

\[\sin2x+2sinx=0 \rightarrow 2sinx.cosx+2sinx=0 \rightarrow 2sinx(cosx+1)=0\] then sinx=0 or cosx+1=0 \[sinx=0 \rightarrow x=k \pi \] \[cosx+1=0 \rightarrow cosx=-1 \rightarrow x=k \pi + \pi\] then the answers are \[\left\{ 0, \pi \right\}\]